OpenStudy (anonymous):

What is the equation of a line that passes through the point (-3, 1) and is parallel to a line with a slope of 2

OpenStudy (abbles):

I'm assuming you need the equation in y-intercept form, which is y = mx + b. To figure that out you must first put it into point-slope form (y - y1 = m(x - x1)) and then isolate the y. Do you know how to get it into point-slope form?

OpenStudy (abbles):

I'll give you a hint. Point-slope form is y - y1 = m(x - x1) where y1 is the y coordinate in (-3, 1) and x1 is the x coordinate. M is the slope.

OpenStudy (anonymous):

igot it thanks

OpenStudy (abbles):

Yep no problem, good luck.